How many lb. of a metal containing 50% silver must be combined with 9 lb. of a metal containing 5% silver to form an alloy containing 23 % silver?
Answer by richwmiller(17219)

0.5*x+0.05*9=0.23(9+x)
0.5*x+0.45=2.07+0.23x
0.5x-0.23x=2.07-0.45
0.27x=1.62
x=6 lbs of 50% is added to 9 lbs of 5% to get 15 lbs at 23%
check
0.5*6+0.05*9=0.23(9+6)
3+0.45=0.23(15)
3.45=3.45
ok